OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) − 4

OpenStudy (dumbcow):

distribute...combine like terms

OpenStudy (anonymous):

did you try this on your own? what was your answer?

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

x=-5/6 sorry haha

OpenStudy (anonymous):

distribute: 3x+3=-2x+1-4 move the -2x over so it becomes 2x 5x+3=+1-4 5x+3=-3 move the 3 over so it becomes -3 5x=-6 get x alone and divide by 5 x=1.2